OpenStudy (anonymous):

Store A: $25 for 45 minutes Store B: $30 for 1 hour Store C: $45 for 1.5 hours Store D: $80 for three 1 hour lessons Julie needs to have swim lessons for her son. There are four different stores that offer swim lessons. Which one is the BEST deal? A) Store A B) Store B C) Store C D) Store D

OpenStudy (monkeys):

D

OpenStudy (igreen):

Divide and see the cost per minute.

OpenStudy (leonardo0430):

D is always the answer monkeh :P

OpenStudy (igreen):

@monkeys Don't give answers.

OpenStudy (monkeys):

oh srry

OpenStudy (igreen):

Store A = $25 for 45 minutes = 45 / 25 = $1.80 a minute Store B = $30 for 60 minutes = 60 / 30 = $2.00 a minute Store C = $45 for 90 minutes = 90 / 45 = $2.00 a minute Store D = $80 for 180 minutes = 180 / 80 = $2.25 a minute So which one is the best deal? @TheUnknownShadow

OpenStudy (igreen):

@monkeys Answer was completely wrong..

OpenStudy (anonymous):

yeah i agree, i went with C